# Hierarchical Augmentation and Distillation for Class Incremental Audio-Visual Video Recognition

### 11 Jan 2024 v1

@@ 1 Introduction
  • The evaluations on four benchmarks demonstrate the superiority of the proposed framework, e.g., obtaining Average Incremental Accuracy / Final Incremental Accuracy of 88.9%/85.1% (87.0%/83.1%), 90.1%/86.6% (89.8%/86.3%), 84.6%/78.0% (84.3%/77.6%) and 78.2%/69.5% (77.6%/69.1%) in AVE 3 phases (6 phases), AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and Kinetics-400 20 phases (40 phases), respectively.
@@ 3 Methods
Given a video $x$, we divide it into $K$ disjoint audio and visual snippet pairs, e.g., $x=\{A_{i},V_{i}\}_{i=1}^{K}$, where $A_{i}$ and $V_{i}$ represent the audio and visual data of the $i$-th video snippet, respectively. The audio-visual model aims to classify the video $x$ by considering all audio snippets $\{A_{i}\}_{i=1}^{K}$ and visual snippets $\{V_{i}\}_{i=1}^{K}$. The audio-visual model $\Phi$ comprises three components: the audio-visual embedding module $\mathcal{E}$, the audio-visual fusion module $\mathcal{F}$, and the classifier module $\mathcal{C}$. The audio-visual embedding module $\mathcal{E}$ employs pre-trained and freezed audio and visual models to extract the low-level modal features $\mathbb{F}=\{f_{i}^{a},f_{i}^{v}\}_{i=1}^{K}$, which consist of audio snippet-level feature $f_{i}^{a}$ and visual snippet-level feature $f_{i}^{v}$. Since storing the videos of past classes demands significant memory, low-level modal features of exemplar videos are retained for knowledge preservation. Specifically, for incremental learning at the $t$-th task, the low-level modal features $\mathbb{F}^{m}$ of the exemplar data $x^{m}$ from the previous $t-1$ tasks are stored in the memory bank $\mathcal{M}_{t-1}$ for knowledge preservation. Moreover, we denote the set of low-level modal features $\mathbb{F}^{c}$ of current data $x^{c}$ in $\mathcal{T}_{t}$ as $\mathcal{T}^{\prime}_{t}$. The audio-visual fusion module $\mathcal{F}$ utilizes the hybrid attention network [55] to perform multi-modal fusion between $\{f_{i}^{a}\}_{i=1}^{K}$ and $\{f_{i}^{v}\}_{i=1}^{K}$, yielding fused audio features $\{h_{i}^{a}\}_{i=1}^{K}$ and visual features $\{h_{i}^{v}\}_{i=1}^{K}$. The video feature $H$ is obtained by fusing the visual and audio features using average pooling,
@@ 3.3.1 Hierarchical Logical Distillation
Unfortunately, the distributions $\mathcal{D}_{1:t-1}$ and $\mathcal{D}_{t}$ are intractable due to the lack of the explicit probability density function. Therefore, we utilize all the given low-level modal features in the memory bank $\mathcal{M}_{t-1}$ and current task $\mathcal{T}^{\prime}_{t}$ to obtain proxy distribution of $\mathcal{D}_{1:t-1}$ and $\mathcal{D}_{t}$. Specifically, the proxy distributions of $\mathcal{D}_{1:t-1}$ and $\mathcal{D}_{t}$ are constructed by the convex hull $\mathcal{H}(\mathcal{M}_{t-1})$ and $\mathcal{H}(\mathcal{T}^{\prime}_{t})$, using all convex combinations of the low-level modal features in the set $\mathcal{M}_{t-1}$ and set $\mathcal{T}^{\prime}_{t}$, respectively:
@@ 4.2 Effect of Hierarchical Augmentation
As the audio-visual embedding module $\mathcal{E}$ is freezed in the audio-visual model $\Phi$, we only concentrate on the audio-visual fusion module $\mathcal{F}$ and classifier module $\mathcal{C}$. We denote the low-level modal feature as $\mathcal{T}$, and $\mathcal{F}(\mathcal{T})$ represents the high-level video feature. We split the neural network $f_{w^{*}}$ into two part: audio-visual fusion module $\mathcal{F}$ and classifier module $\mathcal{C}$ , whose parameters are $w^{*}_{\mathcal{F}}$ and $w^{*}_{\mathcal{C}}$, respectively. With the augmented dataset $\mathcal{T}^{\prime}$, we have:
@@ 5.1 Training Details
Datasets. We conduct class incremental audio-visual video recognition using the AVE, AVK-100, AVK-200, and Kinetics-400 datasets. The AVE dataset [54], derived from AudioSet [71], includes 4,143 videos across 28 categories, divided into 3,339 for training, 402 for validation, and 402 for evaluation. The AVK-100 and AVK-200, on the other hand, are datasets specifically created for class incremental audio-visual video recognition tasks, sourced from the Kinetics-400 without any additional videos. They comprise 100 and 200 distinct classes, respectively, carefully chosen to ensure no class overlap between the two. AVK-100 comprises 59,770 videos from 100 classes, with 35,826 for training, 11,955 for validation and 11,989 for evaluation. AVK-200 contains 114,000 videos and 200 classes, with 68,320 for training, 22,798 for validation and 22,882 for evaluation. Additionally, Kinetics-400 is also used as a dataset for class incremental audio-visual video recognition, abbreviated as AVK-400. It includes 400 classes, with 140,497 videos for training, 46,885 for validation, and 47,045 for testing.
@@
TABLE III: Hierarchy in HAM module on AVE 3 phases.
TABLE IV: Hierarchy in HLD module on AVE 3 phases.
TABLE V: Hierarchy in HCD module on AVE 3 phases.
@@
TABLE IV: Hierarchy in HLD module on AVE 3 phases.
TABLE V: Hierarchy in HCD module on AVE 3 phases.
$AIA=\frac{1}{S}\sum_{i=1}^{S}{IA}_{i},$ (35)
$FIA={IA}_{s}.$ (36)
[…]
TABLE VI: Component analysis of HAD framework on AVE 3 phases.
Method HAM HLD HCD AIA
Baseline 74.8
HAM 87.6
HDM 80.5
78.2
82.2
HAD 88.9
@@ 5.2 Comparison with Existing Methods
In this section, we compare the proposed method with classical exemplar-based methods, such as iCaRL [10], Lucir [35], il2m [11], DER[72], CCIL [13], and CSCCT [42]. These methods use a random sampling strategy to construct the exemplar set for fair comparison. The comparative results are shown in Table II and Figure 4. The term ‘Baseline’ represents using the fine-tune strategy to infer the model on each incoming dataset with the classifier used in our method. Specifically, ‘Baseline’ only concentrates on the audio-visual video recognition of current task by performing classification loss for current task data, and past task data are not available for augmentation or distillation. As the number of tasks increases, the performance of all methods declines overall, which demonstrates the necessity of class incremental audio-visual video recognition task. Compared with the ‘Baseline’, HAD achieves improvements of 14.1%/27.9% (15.5%/28.1%), 31.5%/49.1% (37.9%/57.3%), 43.1%/60.7% (49.0%/57.3%), and 48.5%/58.8% (50.0%/59.2%) about Average Incremental Accuracy / Final Incremental Accuracy metrics in AVE 3 phases (6 phases), AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and AVK-400 20 phases (40 phases), respectively.
Additionally, HAD outperforms existing methods on all datasets, e.g., obtaining Average Incremental Accuracy / Final Incremental Accuracy of 88.9%/85.1% (87.0%/83.1%), 90.1%/86.6% (89.8%/86.3%), 84.6%/78.0% (84.3%/77.6%) and 78.2%/69.5% (77.6%/69.1%) in AVE 3 phases (6 phases), AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and Kinetics-400 20 phases (40 phases), respectively. Moreover, as illustrated in Figure 4, HAD surpasses other methods in nearly all incremental phases, demonstrating its superiority and stability. From Table II, we also observe that existing method ‘il2m’ yields favorable results in large dataset, e.g., 88.4%/84.6% (87.5%/84.3%), 82.9%/75.7% (82.0%/74.9%), and 76.4%/67.2% (75.7%/67.7%) about Average Incremental Accuracy / Final Incremental Accuracy in AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and AVK-400 20 phases (40 phases), respectively. This outcome is attributed to the method’s storage of exemplar data and statistics of old classes. In contrast to ‘il2m’, HAD only stores exemplar data for knowledge preservation. Despite storing fewer exemplars, HAD still outperforms ‘il2m’ across all four datasets. The superior results confirm the effectiveness of the proposed HAD.
@@ 5.3.1 Elements of HAD
We analyze the role of HAM, HDM, HLD, and HCD within the proposed HAD framework based on the setting of the AVE 3 phases with Average Incremental Accuracy (AIA) metric. Table VI shows that using only HAM yields a 12.8% improvement compared with the baseline, indicating that applying hierarchical feature augmentation to exemplar data enhances knowledge preservation. Using HLD and HCD improves the baseline by 5.7% and 3.4%, respectively, demonstrating that both logical distillation and correlative distillation are helpful for model knowledge retention. The comparison of HAD, HLD, and HCD (87.6% v.s 80.5% v.s 78.2%) reveals that the hierarchical augmentation module effectively mitigates catastrophic forgetting in class incremental audio-visual video recognition, with hierarchical logical distillation slightly outperforming hierarchical correlative distillation. By combining HLD and HCD, HDM obtains an average improvement of 7.4%, proving that logical distillation and correlative distillation complement each other. By combining HAM and HDM, HAD improves Average Incremental Accuracy (AIA) by 1.3% and 6.7% for HAM and HDM, respectively. The superior performance indicates that data knowledge preservation and model knowledge preservation are essential for class incremental audio-visual video recognition.
@@ 5.3.2 Effect of Hierarchical Structure
We analyze the necessity of hierarchical structure in feature augmentation, i.e., logical distillation and correlative distillation with Average Incremental Accuracy (AIA) metric, and summarize the results in Table V, Table V, and Table V, respectively.
Table V demonstrates that employing the low-level modal augmentation (LMA) and high-level video augmentation (HVA) both result in higher performance compared with the baseline. For example, Average Incremental Accuracy (AIA) is improved from 74.8% to 86.4%/87.0% for LMA/HVA, indicating that both low-level feature augmentation and high-level feature augmentation effectively maintain knowledge of previous classes. By combining LMA and HVA, HAM obtains an Average Incremental Accuracy (AIA) of 87.6%, suggesting that low-level modal augmentation and high-level video augmentation reinforce each other. Consequently, it is necessary to consider both low-level and high-level feature augmentations.
Table V shows that using the video-level logical distillations (SLD) and distribution-level logical distillations (DLD) yields improvements of 5.5% and 4.9% compared with the baseline, respectively, illustrating that both video-level logical distillation and distribution-level logical distillation effectively preserve model knowledge. Combining SLD and DLD achieves the best Average Incremental Accuracy (AIA) of 80.5%, verifying the necessity of hierarchical logical distillation.
Table V indicates that employing the snippet-level correlative distillation (SCD) and video-level correlative distillation (VCD) leads to improvements of 2.1% and 3.1% over the baseline, respectively, demonstrating the effectiveness of snippet-level correlative distillation and video-level correlative distillation. HCD performs better by combining SCD and VCD, supporting the rationale behind hierarchical distillation.
From Table V, Table V, and Table V, we can conclude that it is necessary to consider the hierarchical structure of the model and data to preserve data knowledge and model knowledge for class incremental audio-visual video recognition.
@@ 5.3.6 Hyperparameter Analysis
5.3.8 Compared with video class incremental learning method
To verify the effectiveness of our method and the validity of the proposed Class Incremental Audio-Visual Video Recognition paradigm, we compare our approach with the state-of-the-art video class incremental learning method PIVOT [76]. Note that PIVOT focuses solely on video, while our method targets multi-modal incremental learning for both audio and video. For making a fair comparison, we unify the modality to conduct experiments and report the results with Final Incremental Accuracy (FIA) metric shown in Table VII. We conduct class incremental learning for visual modality (HAD-V) and multi-modality (HAD) on the Kinetics-400 dataset by equally dividing its 400 classes into either 10 or 20 distinct tasks. Comparing HAD with the state-of-the-art video class incremental learning method PIVOT [76], it can be observed that the performance PIVOT exceed that of HAD-V by 0.3% and 0.7% in 10 Tasks and 20 Tasks, respectively. The reason is that PIVOT employs the large-scale transformers VIT-B/32 [77] as the backbone, while the proposed method uses the smaller ResNet-152 [7] and 3D ResNet [74] as the backbone. However, HAD considering both visual and audio information obtains a higher performance than PIVOT, e.g., surpassing PIVOT by 9.3% and 9.6% in 10 Tasks and 20 Tasks, respectively. The superior performance demonstrates the effectiveness and rationality of considering the visual and audio information for Class Incremental Audio-Visual Video Recognition.
TABLE VII: Final Incremental Accuracy (FIA) of class incremental learning in AVK-400 by equally dividing all the classes into 10 or 20 tasks.
Methods Visual Audio AA (10 Tasks) AA (20 Tasks)
PIVOT [76] 58.6 57.5
HAD-V (Ours) 58.3 56.8
HAD (Ours) 67.9 67.1

### 10 Apr 2024 v2

@@ 1 Introduction
  • The evaluations on four benchmarks demonstrate the superiority of the proposed framework, e.g., obtaining Average Incremental Accuracy / Final Incremental Accuracy of 88.9%/85.1% (87.0%/83.1%), 90.1%/86.6% (89.8%/86.3%), 84.6%/78.0% (84.3%/77.6%) and 78.2%/69.5% (77.6%/69.1%) in AVE 3 phases (6 phases), AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and AVK-400 20 phases (40 phases), respectively.
@@ 3 Methods
Given a video $x$, we divide it into $K$ disjoint audio and visual snippet pairs, e.g., $x=\{A_{i},V_{i}\}_{i=1}^{K}$, where $A_{i}$ and $V_{i}$ represent the audio and visual data of the $i$-th video snippet, respectively. The audio-visual model aims to classify the video $x$ by considering all audio snippets $\{A_{i}\}_{i=1}^{K}$ and visual snippets $\{V_{i}\}_{i=1}^{K}$. The audio-visual model $\Phi$ comprises three components: the audio-visual embedding module $\mathcal{E}$, the audio-visual fusion module $\mathcal{F}$, and the classifier module $\mathcal{C}$. The audio-visual embedding module $\mathcal{E}$ employs pre-trained and frozen audio and visual models to extract the low-level modal features $\mathbb{F}=\{f_{i}^{a},f_{i}^{v}\}_{i=1}^{K}$, which consist of audio snippet-level feature $f_{i}^{a}$ and visual snippet-level feature $f_{i}^{v}$. Therefore, the primary emphasis of CIAVVR is on the learning and catastrophic forgetting of the audio-visual fusion module and classifier module, rather than on the catastrophic forgetting of the frozen pre-trained audio-visual embedding module. Since storing the videos of past classes demands significant memory, low-level modal features of exemplar videos are retained for knowledge preservation. Specifically, for incremental learning at the $t$-th task, the low-level modal features $\mathbb{F}^{m}$ of the exemplar data $x^{m}$ from the previous $t-1$ tasks are stored in the memory bank $\mathcal{M}_{t-1}$ for knowledge preservation. Moreover, we denote the set of low-level modal features $\mathbb{F}^{c}$ of current data $x^{c}$ in $\mathcal{T}_{t}$ as $\mathcal{T}^{\prime}_{t}$. The audio-visual fusion module $\mathcal{F}$ utilizes the hybrid attention network [55] to perform multi-modal fusion between $\{f_{i}^{a}\}_{i=1}^{K}$ and $\{f_{i}^{v}\}_{i=1}^{K}$, yielding fused audio features $\{h_{i}^{a}\}_{i=1}^{K}$ and visual features $\{h_{i}^{v}\}_{i=1}^{K}$. The video feature $H$ is obtained by fusing the visual and audio features using average pooling,
@@ 3.3.1 Hierarchical Logical Distillation
Therefore, we utilize all the given low-level modal features in the memory bank $\mathcal{M}_{t-1}$ and current task $\mathcal{T}^{\prime}_{t}$ to obtain proxy distribution of $\mathcal{D}_{1:t-1}$ and $\mathcal{D}_{t}$. Specifically, the proxy distributions of $\mathcal{D}_{1:t-1}$ and $\mathcal{D}_{t}$ are constructed by the convex hull $\mathcal{H}(\mathcal{M}_{t-1})$ and $\mathcal{H}(\mathcal{T}^{\prime}_{t})$, using all convex combinations of the low-level modal features in the set $\mathcal{M}_{t-1}$ and set $\mathcal{T}^{\prime}_{t}$, respectively:
@@ 4.2 Effect of Hierarchical Augmentation
As the audio-visual embedding module $\mathcal{E}$ is frozen in the audio-visual model $\Phi$, we only concentrate on the audio-visual fusion module $\mathcal{F}$ and classifier module $\mathcal{C}$. We denote the low-level modal feature as $\mathcal{T}$, and $\mathcal{F}(\mathcal{T})$ represents the high-level video feature. We split the neural network $f_{w^{*}}$ into two part: audio-visual fusion module $\mathcal{F}$ and classifier module $\mathcal{C}$ , whose parameters are $w^{*}_{\mathcal{F}}$ and $w^{*}_{\mathcal{C}}$, respectively. With the augmented dataset $\mathcal{T}^{\prime}$, we have:
@@ 5.1 Training Details
Datasets. We use the AVE, AVK-100, AVK-200, and AVK-400 datasets for class-incremental audio-visual video recognition. The AVE dataset [54], derived from AudioSet [71], includes 4,143 videos across 28 categories, with 3,339 for training, 402 for validation, and 402 for evaluation. On the other hand, AVK-100, AVK-200, and AVK-400 are datasets specifically created for class-incremental audio-visual video recognition tasks, sourced from Kinetics-400 [72] without any additional videos. Due to some videos in Kinetics-400 having invalid YouTube download links and issues with extracting audio-visual features, we constructed the AVK-100, AVK-200, and AVK-400 datasets based on Kinetics-400 and divided them into training, validation, and evaluation sets ourselves. AVK-100 contains 59,770 videos from 100 categories, with 35,826 for training, 11,955 for validation, and 11,989 for evaluation. AVK-200 contains 114,000 videos from 200 categories, with 68,320 for training, 22,798 for validation, and 22,882 for evaluation. AVK-400 includes 234,427 videos from 400 categories, with 140,497 for training, 46,885 for validation, and 47,045 for evaluation.
@@
$AIA=\frac{1}{S}\sum_{i=1}^{S}{IA}_{i},$ (35)
$FIA={IA}_{s}.$ (36)
Implementation Details
TABLE III: Component analysis of HAD framework on AVE 3 phases.
Method HAM HLD HCD AIA
Baseline 74.8
HAM 87.6
HDM 80.5
78.2
82.2
HAD 88.9
[…]
TABLE IV: Hierarchy in HAM module on AVE 3 phases.
TABLE V: Hierarchy in HLD module on AVE 3 phases.
TABLE VI: Hierarchy in HCD module on AVE 3 phases.
@@
TABLE V: Hierarchy in HLD module on AVE 3 phases.
TABLE VI: Hierarchy in HCD module on AVE 3 phases.
@@ 5.2 Comparison with Existing Methods
In this section, we compare the proposed method with classical exemplar-based methods, such as iCaRL [10], Lucir [35], il2m [11], DER[73], CCIL [13], and CSCCT [42]. To ensure a fair comparison, all these methods are evaluated on the same dataset division and leverage both audio and visual features of the videos. Each method is re-implemented using the same audio and visual features generated by the pre-trained audio-visual embedding module. Moreover, a random sampling strategy is employed to construct the exemplar set for each method. The comparative results are shown in Table II and Figure 4. The term ‘Baseline’ represents using the fine-tune strategy to infer the model on each incoming dataset with the classifier used in our method. Specifically, ‘Baseline’ only concentrates on the audio-visual video recognition of current task by performing classification loss for current task data, and past task data are not available for augmentation or distillation. As the number of tasks increases, the performance of all methods declines overall, which demonstrates the necessity of class incremental audio-visual video recognition task. Compared with the ‘Baseline’, HAD achieves improvements of 14.1%/27.9% (15.5%/28.1%), 31.5%/49.1% (37.9%/57.3%), 43.1%/60.7% (49.0%/57.3%), and 48.5%/58.8% (50.0%/59.2%) about Average Incremental Accuracy / Final Incremental Accuracy metrics in AVE 3 phases (6 phases), AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and AVK-400 20 phases (40 phases), respectively.
Additionally, HAD outperforms existing methods on all datasets, e.g., obtaining Average Incremental Accuracy / Final Incremental Accuracy of 88.9%/85.1% (87.0%/83.1%), 90.1%/86.6% (89.8%/86.3%), 84.6%/78.0% (84.3%/77.6%) and 78.2%/69.5% (77.6%/69.1%) in AVE 3 phases (6 phases), AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and AVK-400 20 phases (40 phases), respectively. Moreover, as illustrated in Figure 4, HAD surpasses other methods in nearly all incremental phases, demonstrating its superiority and stability. From Table II, we also observe that existing method ‘il2m’ yields favorable results in large dataset, e.g., 88.4%/84.6% (87.5%/84.3%), 82.9%/75.7% (82.0%/74.9%), and 76.4%/67.2% (75.7%/67.7%) about Average Incremental Accuracy / Final Incremental Accuracy in AVK-100 5 phases (10 phases), AVK-200 10 phases (20 phases) and AVK-400 20 phases (40 phases), respectively. This outcome is attributed to the method’s storage of exemplar data and statistics of old classes. In contrast to ‘il2m’, HAD only stores exemplar data for knowledge preservation. Despite storing fewer exemplars, HAD still outperforms ‘il2m’ across all four datasets. The superior results confirm the effectiveness of the proposed HAD.
@@ 5.3.1 Elements of HAD
We analyze the role of HAM, HDM, HLD, and HCD within the proposed HAD framework based on the setting of the AVE 3 phases with Average Incremental Accuracy (AIA) metric. Table III shows that using only HAM yields a 12.8% improvement compared with the baseline, indicating that applying hierarchical feature augmentation to exemplar data enhances knowledge preservation. Using HLD and HCD improves the baseline by 5.7% and 3.4%, respectively, demonstrating that both logical distillation and correlative distillation are helpful for model knowledge retention. The comparison of HAD, HLD, and HCD (87.6% v.s 80.5% v.s 78.2%) reveals that the hierarchical augmentation module effectively mitigates catastrophic forgetting in class incremental audio-visual video recognition, with hierarchical logical distillation slightly outperforming hierarchical correlative distillation. By combining HLD and HCD, HDM obtains an average improvement of 7.4%, proving that logical distillation and correlative distillation complement each other. By combining HAM and HDM, HAD improves Average Incremental Accuracy (AIA) by 1.3% and 6.7% for HAM and HDM, respectively. The superior performance indicates that data knowledge preservation and model knowledge preservation are essential for class incremental audio-visual video recognition.
@@ 5.3.2 Effect of Hierarchical Structure
We analyze the necessity of hierarchical structure in feature augmentation, i.e., logical distillation and correlative distillation with Average Incremental Accuracy (AIA) metric, and summarize the results in Table VI, Table VI, and Table VI, respectively.
Table VI demonstrates that employing the low-level modal augmentation (LMA) and high-level video augmentation (HVA) both result in higher performance compared with the baseline. For example, Average Incremental Accuracy (AIA) is improved from 74.8% to 86.4%/87.0% for LMA/HVA, indicating that both low-level feature augmentation and high-level feature augmentation effectively maintain knowledge of previous classes. By combining LMA and HVA, HAM obtains an Average Incremental Accuracy (AIA) of 87.6%, suggesting that low-level modal augmentation and high-level video augmentation reinforce each other. Consequently, it is necessary to consider both low-level and high-level feature augmentations.
Table VI shows that using the video-level logical distillations (SLD) and distribution-level logical distillations (DLD) yields improvements of 5.5% and 4.9% compared with the baseline, respectively, illustrating that both video-level logical distillation and distribution-level logical distillation effectively preserve model knowledge. Combining SLD and DLD achieves the best Average Incremental Accuracy (AIA) of 80.5%, verifying the necessity of hierarchical logical distillation.
Table VI indicates that employing the snippet-level correlative distillation (SCD) and video-level correlative distillation (VCD) leads to improvements of 2.1% and 3.1% over the baseline, respectively, demonstrating the effectiveness of snippet-level correlative distillation and video-level correlative distillation. HCD performs better by combining SCD and VCD, supporting the rationale behind hierarchical distillation.
From Table VI, Table VI, and Table VI, we can conclude that it is necessary to consider the hierarchical structure of the model and data to preserve data knowledge and model knowledge for class incremental audio-visual video recognition.
@@ 5.3.6 Hyperparameter Analysis
5.3.8 Analysis of out-of-distribution dataset
In the main experiments, the AVE dataset is a subset of the Audio-set, which is a pre-trained dataset for the audio feature extraction network VGGish. AVK100, AVK200, and AVK400 are subsets of the Kinetics-400 dataset for the 3D ResNet visual feature extraction network. Therefore, we can obtain effective feature representations from these datasets to provide good inputs for subsequent audio-visual fusion and classification modules. However, for audio-visual incremental learning tasks, the dataset may not always align with the pre-trained model’s dataset, potentially resulting in suboptimal or corrupted features. To assess our model’s generalization capabilities for out-of-distribution datasets, we introduce a new dataset called Extra-AVK. This dataset consists of 200 classes selected from Kinetics-600, which are distinct from the Kinetics-400 classes. We extract 200 videos per class and split them into training, validation, and evaluation sets with a 6:2:2 ratio, resulting in 120 training, 40 validation, and 40 test videos per class. Unlike the other datasets, the Extra-AVK dataset does not have pre-trained information for the audio, 2D visual, and 3D visual networks. We use 100 classes from Extra-AVK as the initial base task and divide the remaining 100 classes into 10 or 20 incremental tasks (phases), with each task containing 10 or 5 classes, respectively.
From the Table VII, it’s evident that our method HAD outperforms other approaches on the Extra-AVK dataset, which is completely new to the audio, 2D visual, and 3D visual networks. For example, HAD achieves the best Average Incremental Accuracy / Final Incremental Accuracy of 58.6%/45.5% (57.0%/44.6%) on Extra-AVK with 10 phases (20 phases), underscoring our method’s effectiveness and generalization capability. Moreover, we observe a significant decline in the performance of Joint-training and Baseline on Extra-AVK compared to their performance on AVK-200 in Table II. Specifically, Joint-training on Extra-AVK sees a 26.0% drop compared to AVK-200. The Baseline method on Extra-AVK with 10 phases (20 phases) shows a decrease in Average Incremental Accuracy / Final Incremental Accuracy by 27.9%/13.4% (25.9%/18.1%) compared to AVK-200 with 10 phases (20 phases). This decline indicates that the pre-trained audio-visual embedding module lacks sufficient prior knowledge for the Extra-AVK dataset, leading to inferior feature extraction and, consequently, poorer audio-visual classification results. This suggests that for future audio-visual incremental learning tasks, employing more advanced audio-visual embedding module could enhance performance.
TABLE VII: Average Incremental Accuracy (AIA) / Final Incremental Accuracy (FIA) in Extra-AVK with 10/20 phases.
Methods Extra-AVK
No. of incremental tasks 10 20
Baseline 13.6 / 3.9 9.4 / 2.2
iCaRL [10] 51.7 / 37.1 47.7 / 28.6
Lucir [35] 33.0 / 22.8 31.6 / 22.0
il2m [11] 56.0 / 42.5 53.7 / 39.8
DER [73] 43.1 / 25.4 38.3 / 25.0
CCIL [13] 36.7 / 29.4 33.6 / 26.6
CSCCT [42] 55.4 / 42.5 51.6 / 37.8
HAD(Ours) 58.6 / 45.5 57.0 / 44.6
Joint-training 62.8 62.8